I have a triangle and it ask find the measurement of A the measurement are 130 degrees then 2x degrees then 3x degrees how do I solve
Answer by mananth(16946)

The sum of angles of triangle = 180 degrees
2x+3x + 130=180
5x=180-130
5x=50
x=10
2x = 2*10 = 20 degrees
3x=3*10=30 degrees